InsatiableSuffering:

Can somebody explain how to integrate sin(2t^4)t^7dt? Do you have to use u substitution?

InsatiableSuffering:

@justjm do you think you know

justjm:

is it \(\int_{ }^{ }t^{7}\sin\left(t^{4}\right)dt\) because then you need to do integration by parts AND u-substitution

InsatiableSuffering:

Yeah, it's that integral. So what would u be set equal to in this case? Would u be equal to 2t^4 or t^4 or t^7?

justjm:

So since the t^7 is so big...you would end up doing integration by parts 8 times!!!! but instead, you can apply u-sub first on the t^4 first, and get ride of the high degree. First do a u-sub and let u=t^4. Then you can cancel out most of the t^7 as you will notice. Then you do integration by parts

justjm:

Hold on let me try doing it

justjm:

\(\int_{ }^{ }t^{7}\sin\left(t^{4}\right)dt\) let u=t^4 \(\frac{1}{4}\int_{ }^{ }t^{7}\sin\left(u\right)\frac{du}{t^{3}}\) \(\frac{1}{4}\int_{ }^{ }t^{4}\sin\left(u\right)du\) \(\frac{1}{4}\int_{ }^{ }u\sin\left(u\right)du\) now use integration by parts let u=u, so du=du let dv=sinudu, so v=-cosu Formula for integration by parts: \(uv-\int_{ }^{ }v'du\) \(\frac{1}{4}\left(-u\cos u-\int_{ }^{ }-\cos udu\right)\) \(\frac{1}{4}\left(-\cos u+\sin u+c\right)\) YOu can plug u=t^4 back in, but all of the u's might've made it a bit confusing oof

InsatiableSuffering:

Oh my bad

InsatiableSuffering:

@justjm I looked at the integral incorrectly

InsatiableSuffering:

It was that, but with the sin, it was actually 2*t^4

justjm:

Oh it shouldn't be much an issue, you'd still do the same thing. Do u-sub first then the integration by parts.

InsatiableSuffering:

Thank you so much, that helped a lot.